OpenStudy (anonymous):

solve by factoring 4x^2+19x=-21

OpenStudy (anonymous):

4x^2+19x=-21 4x^2+19x+21=0 divide by 4 x^2 +(19/4)x+(21/4)=0 (x+3)(x+ 7/4)=0 ans